# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

### v1

@@ 3.1 End-to-end Latency of RAG Pipelines
Figure 3: Latency breakdown of six RAG pipelines on NQ dataset [54].
@@
Figure 4: The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval. The dotted line indicates the memory capacity of a RTX4090 GPU, which is a common used GPU for local deployment.
@@
We begin by analyzing the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios. In the first, the LLM runs on the GPU while retrieval is offloaded to the CPU, minimizing GPU memory usage. In the second, both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency. If the GPU’s memory is insufficient, then CPU offloading is necessary.
@@
We set the nprobe to 512, a common used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency. Figure 3 shows the breakdown of end-to-end latency into LLM computation and retrieval. The results are the average of 512 random data from the NQ dataset [54]. We observe that CPU-based retrieval (first scenario) is the primary bottleneck, consuming about 56% of the total execution time on average, while GPU-based retrieval (second scenario) accounts for just 12%. On average, GPU retrieval is 12.5x faster than CPU retrieval, reducing overall latency by approximately 2.2x. Thus, accelerating retrieval on GPU is crucial for improving end-to-end latency.
@@
However, GPU acceleration comes with a significant memory cost. Figure 4 shows the memory requirements for both GPU and CPU. As Figure 4 shows, putting both LLM weights and the retrieval index on the GPU requires around 77GB of memory, which exceeds the capacity of consumer GPUs like the RTX 4090 with 24GB. Thus, GPU acceleration on retrieval is often not feasible when running on a lower-end GPU or indexing with a large datastore. Note that the memory footprint of the LLM and the index can be even larger and further exacerbate the memory constraint.
@@
In the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
3.2 Opportunities of GPU-accelerated Retrieval with Limited Memory
A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the target search clusters (§2.2). While such a method enables faster search on the GPU, the bottleneck of retrieval is shifted to the data fetching.
To examine the performance, we reconstruct the IVF index in PyTorch [8], leveraging its high-performance GPU operator, and implement such an on-demand fetching retrieval system. We measure the retrieval latency with an RTX 4090 GPU and compare it with the baseline of CPU-based retrieval. Figure 5 shows the breakdown of retrieval time into data fetch and search time. We show the performance for different amounts of data fetched (indicated by the nprobe parameter). We observe that the fetch time dominates the latency due to limited PCIe bandwidth between the CPU and GPU (32 GB/s). Despite the fact that GPU retrieval is significantly faster than CPU retrieval, due to the fetch overhead, the effective latency is slower than CPU retrieval (around 3% slow down on average across nprobe’s). Therefore, we must hide the data fetching latency in order to achieve a meaningful speedup to following such approach.
Hiding the data fetch cost. Hiding the data fetch cost implies that CPU-to-GPU data fetch must be done before the retrieval operation is performed. However, predicting which data to retrieve in advance is challenging since the exact data requirements only become clear at retrieval time once we have the query produced by the pre-retrieval process (§2.1).
Fortunately, as we show next, the original form of a query before pre-retrieval can serve as a valuable hint for the clusters required to be retrieved at the end of the pre-retrieval stage. This hint can guide us in predicting the clusters that should be fetched, and the data transfer can be initiated in the background along with pre-retrieval to hide its latency.
@@ 3.2 Opportunities of GPU-accelerated Retrieval with Limited Memory
Figure 5: Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
@@
[27] [58] [58] [78] [43] [11]
NQ [54] 75.7% 66.2% 93.4% 85.7% 80.3% 100.0%
HotpotQA [91] 78.6% 66.6% 86.9% 91.2% 82.2% 100.0%
TriviaQA [47] 76.8% 65.3% 86.8% 87.5% 82.9% 100.0%
Table 1: IVF cluster overlapping rate between the input and output of the pre-retrieval generation. The nprobe is set to 512. Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
3.3 Overlapping of IVF Clusters
@@
Similarity of queries at different stages. During the pre-retrieval process of RAG pipelines (e.g., query transformation [61, 37, 27, 102, 93, 73, 72, 100, 58]), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval. This process often rewrites the query into a different format [27] or simplifies its complexities [58], which intuitively preserves the query’s core semantic content. Hence, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar. This similarity, in turn, suggests that the IVF clusters to which these queries would be assigned will overlap. Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting subsequent memory accesses.
Prediction coverage. To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ [54], HotpotQA [91], and TriviaQA [47]) and six RAG pipelines. Table 1 shows the coverage when we prefetch $512$ clusters. From the table, we observe that IVF cluster overlap rates are consistently high across a range of datasets and pipelines. For instance, even the lowest reported values remain above 65% (for SubQ).
@@
4 Design of TeleRAG
Figure 6: (a) The overview of TeleRAG’s lookahead retrieval and comparison to the baseline CPU-offloaded retrieval. (b) System design of TeleRAG. It first loads the entire index data into CPU memory and identifies the clusters ($C_{\text{in}}$) for the query before the transformation ($q_{\text{in}}$). It then transfers the data of $C_{\text{in}}$ to GPU memory while running the pre-retrieval stage and generating the retrieval query $q_{\text{out}}$. During the search, the overlapped clusters are searched on GPU, and the missed clusters are searched on CPU. The retrieval results are then merged and passed to the next stage.
@@ 4 Design of TeleRAG
Based on the high IVF cluster overlapping observation presented in § 3.3, we present TeleRAG, an efficient inference system for RAG that incorporates the idea of lookahead retrieval, which predicts and prefetches likely IVF cluster data to GPU and significantly accelerates the retrieval process with minimal GPU memory requirement. In this section, we describe our design overview (§4.1), an analytical model on finding the optimal prefetch data amount (§4.3), and the implementation details (§4.4).
4.1 Overview
@@
Figure 6 presents the overview of TeleRAG with lookahead retrieval. In Figure 6, the input to the pre-retrieval stage is defined as $q_{\text{in}}$, and the output is denoted as $q_{\text{out}}$. In the subsequent retrieval stage, $q_{\text{out}}$ serves as the input query to perform the retrieval. We highlight the IVF clusters corresponding to $q_{\text{in}}$ with a red background ($C_{\text{in}}$), and those corresponding to $q_{\text{out}}$ with a blue background ($C_{\text{out}}$). Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$3.3), marked with a purple background ($C_{\text{overlap}}$).
Given this, the lookahead retrieval technique in TeleRAG operates in three key steps. First, $C_{\text{in}}$ is computed, and their data is transferred to GPU memory during pre-retrieval generation, beginning with those clusters whose centroids are closest to $q_{\text{in}}$. This is possible because GPUs are equipped with Direct Memory Access (DMA) engines, which can handle memory transfers concurrently with compute operations. Additionally, LLM generation is typically a compute-intensive task that involves minimal data transfer between the device and host memory. Hence, lookahead retrieval can run seamlessly alongside LLM generation.
Second, at retrieval stage, the GPU performs a fast similarity search on the accurately predicted clusters ($C_{\text{overlap}}$) after determining $C_{\text{out}}$. Since similarity search essentially involves dot-product computations with high dimension embeddings, this is highly efficient on the GPU.
Third, simultaneously to the second step, the CPU processes the similarity search for remaining clusters of $C_{\text{out}}$ that were not prefetched and is stored on CPU memory. Although CPUs are generally slower at performing similarity search, lookahead retrieval significantly reduces the workload assigned to the CPU, thereby accelerating the overall retrieval process.
Additionally, TeleRAG optimizes the sorting of vector distances, which is required for the $k$-argmin operation, the final step of the IVF search where the top-$k$ indices are identified. Since each cluster could have more than thousands of embedding vectors, sorting is much faster on the GPU than on the CPU [9]. In order to perform the sorting on GPU, TeleRAG passes the CPU computed distances of $C_{\text{miss}}$ to GPU first and then perform the global sorting on the combined distances of $C_{\text{overlap}}$ and $C_{\text{miss}}$. Unlike the original vector data, the distance values (such as L2-norm or inner product [46]) are scalars per data point and thus, are much smaller in size. Hence, transferring them to the GPU incurs minimal overhead, allowing the GPU to achieve end-to-end speedup on the sorting process. This GPU-based sorting optimization is typically absent in CPU-based retrieval implementations [25] as they assume there is no GPU in the systems.
In summary, TeleRAG optimizes retrieval by leveraging lookahead retrieval to hide the data movement time for prefetching, GPU computation for efficient similarity search and sorting, and CPU processing for non-overlapping clusters, significantly accelerating the retrieval process.
4.2 Prefetching Design
Prefetch target. In the early design of TeleRAG, we use number of clusters as the prefetch target as it aligns with the nprobe setting of the IVF retrieval process. With this target, Despite we can configure the number clusters to prefetch to adapt to each pipeline’s characteristics, we found this target brings an unstable inference performance. It’s because the size of each cluster is often heavily skewed [59, 44], making data loading time for some clusters much longer than the others. To solve this, we use number of bytes, $b_{p}$, as the prefetch target. With $b_{p}$ as the prefeth target, it provides an upper bound on prefetching time at $\frac{b_{p}}{B}$, where $B$ stands for the CPU-GPU memory bandwidth.
To perform prefetching, we first sort the clusters by the distance between their centroids and the query embedding vector. Starting with the nearest cluster, we include clusters one by one until adding another would exceed the target $b_{p}$. We do not split clusters to completely fill $b_{p}$, ensuring a clear separation between GPU and CPU searches.
Prefetch for Multi-Round RAG. As discussed in §2.1, RAG pipelines often run through multiple rounds. Because each round still processes for the same input query, the IVF clusters used in retrieval are highly similar across rounds. In TeleRAG, we therefore perform a full prefetch only for the first round, then fetch only the additional IVF clusters needed for subsequent rounds, until reaching the memory budget. We currently flush the prefetched data for each new query, assuming no correlation across queries. We leave more advanced memory management and identification of hot clusters between queries for future work.
4.3 Finding the Optimal Prefetch Amount
@@
A key challenge in TeleRAG is to balance the benefit of reducing retrieval latency by prefetching data against the overhead of CPU-GPU transfers. Prefetching more clusters reduces the subsequent retrieval time but can also extend the transfer phase; if it extends beyond the LLM generation window, we lose the advantage of overlap and potentially introduce additional delay. Nevertheless, even additional delay in transferring data can be worthwhile if it substantially reduces retrieval latency.
Mathematical model. We develop a mathematical model to guide the choice of the optimal amount of data to prefetch. Here, we follow the notations from §4.2 where $b_{p}$ is bytes to prefetch and $B$ is the CPU-GPU bandwidth. Optimal amount of data to prefetch is denoted as $b_{p}^{*}$.
To start, we let $t_{1}$ represent the combined time of prefetching and pre-retrieval LLM generation, and $t_{2}$ represents the retrieval time. We have:
$t_{1}$ $=\max(t_{\text{LLM}},t_{p}),$ (1)
$t_{2}$ $=\max(t_{c},t_{g}),$ (2)
where $t_{\text{LLM}}$ is the time for LLM generation, $t_{p}$ is prefetching time, $t_{c}$ is the CPU retrieval time, and $t_{g}$ is the GPU retrieval time. The objective is to minimize $t_{1}+t_{2}$.
@@
$t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$ (3)
@@
As shown in §3.1, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$). Thus, we assume $t_{c}\gg t_{g}$. As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed [40], such that:
$t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$ (4)
@@
    $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$ (6)
@@
    $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$ (7)
@@
4.4 Implementation
TeleRAG is implemented in Python and leverages PyTorch’s [8] operator ecosystem for efficient computation. The datastore index is initially constructed using FAISS [25], and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors. These tensors are stored on disk to eliminate the need for repeated conversions.

### v2

@@ 3.1. End-to-end Latency of RAG Pipelines
Figure 3. Latency breakdown of six RAG pipelines on NQ dataset (Kwiatkowski et al., 2019) with one H100 GPU. nprobe is 256.
@@
Figure 4. The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval. The dotted line indicates the memory capacity of a RTX4090 GPU, which is a common used GPU for local deployment.
@@
We first analyze the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios: (1) the LLM runs on the GPU, while retrieval is offloaded to the CPU, minimizing GPU memory usage, and (2) both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency. If the GPU’s memory is insufficient, CPU offloading is necessary.
@@
We set the nprobe to 256, a commonly used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency. Figure 3 shows the breakdown of end-to-end latency into LLM inference part and retrieval part, based on 1024 randomly sampled queries from the NQ dataset (Kwiatkowski et al., 2019). We observe that, in the CPU-based retrieval baseline (first scenario), the retrieval phase dominates the latency, consuming 41.1% and 60.5% of total latency for batch sizes 1 and 4, respectively. In contrast, GPU-accelerated retrieval in the second scenario dramatically reduces this bottleneck, accounting for only 10.5% and 28.3% of the latency in comparable configurations. On average, GPU retrieval is $5.96\times$ (batch size 1) or $3.87\times$ (batch size 4) faster than CPU retrieval, reducing overall latency by approximately $1.5\times$ or $1.8\times$, respectively. Thus, accelerating retrieval on the GPU is crucial for improving end-to-end latency.
@@
However, GPU acceleration comes with a significant memory cost. Figure 4 shows the memory requirements for both GPU and CPU. As Figure 4 shows, putting both LLM weights and the retrieval index on the GPU requires around 77 GB of memory, which exceeds the capacity of consumer GPUs like the RTX 4090 with 24 GB. Thus, GPU acceleration on retrieval is often not feasible when running on a lower-end GPU or indexing with a large datastore.
Even on GPUs capable of holding the entire index in memory (e.g., H100), the index’s memory footprint limits throughput in data center serving scenario. Batched inference maximizes throughput, but it is bottlenecked by the KV cache capacity for LLM serving (Kwon et al., 2023). For the case of the example above, storing the index on the GPU leaves minimal memory for the KV cache (e.g., ~3 GB, or ~20k tokens worth), restricting batch size, while offloading the index to the CPU significantly increases the available memory (e.g., to ¿500k tokens worth). This issue is exacerbated since RAG typically involves long contexts (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024) and there is a trend of growing model and index sizes (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]), increasing memory pressure and necessitating offloading. However, CPU offloading is suboptimal for latency-sensitive applications as discussed above.
@@
Therefore, in the rest of this section, we try to answer the following question: Is it possible to achieve the latency of GPU-based retrieval while using much less GPU memory?
@@
Figure 5. Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
3.2. Opportunities of GPU-accelerated Retrieval with Limited Memory
A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the search space (§2.2). While this method enables faster searching on the GPU, data fetching becomes the bottleneck.
In Figure 5, we compare the latency of the on-demand fetching system against CPU retrieval on an RTX 4090 GPU. We show three different nprobe values that determine the amount of data fetched. Overall, fetch time dominates latency due to the limited PCIe bandwidth between the CPU and GPU (32 GB/s). Although the GPU search is substantially faster than the CPU counterpart, the fetch overhead results in a higher end-to-end latency, which is about $3\%$ slower on average across nprobe values. Thus, to realize any meaningful speedup with this approach, the data fetching latency must be effectively hidden.
To hide data fetch costs, CPU-to-GPU transfer must precede retrieval, which requires predicting which data will be accessed. Fortunately, as we show next, the modern RAG pipelines offer a valuable hint as a query at the previous step.
@@ 3.2. Opportunities of GPU-accelerated Retrieval with Limited Memory
(Gao et al., 2023a) (Liu, 2022) (Liu, 2022) (Shao et al., 2023) (Jiang et al., 2023c) (Asai et al., 2023)
NQ (Kwiatkowski et al., 2019) 73.1% 63.2% 91.5% 83.8% 79.1% 100.0%
HotpotQA (Yang et al., 2018) 75.3% 62.5% 89.6% 89.4% 80.2% 100.0%
TriviaQA (Joshi et al., 2017) 73.1% 61.6% 86.2% 86.1% 81.7% 100.0%
Table 1. IVF cluster overlapping rate between the input and output of the pre-retrieval generation. The nprobe is set to 512. Since Self-RAG does not incorporate query transform, its coverage is always $100\%$.
3.3. Overlapping of IVF Clusters
@@
Similarity of queries at different stages. During the pre-retrieval process of RAG pipelines (e.g., query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022)), an LLM call is issued to refine an initial user query $q_{\text{in}}$ into a transformed query $q_{\text{out}}$, which is then used for retrieval. This process often rewrites the query into a different format (Gao et al., 2023a) or simplifies its complexities (Liu, 2022), which intuitively preserves the query’s core semantic content. Hence, the embedding vectors of $q_{\text{in}}$ and $q_{\text{out}}$ are likely to be similar. This similarity, in turn, suggests that the IVF clusters to which these queries would be assigned will overlap. Therefore, $q_{\text{in}}$ can serve as a valuable hint for predicting subsequent memory accesses.
Prediction coverage. To verify this hypothesis, we evaluate the average cluster coverage rate between prefetched clusters and clusters used for retrieval in three popular question-answer datasets (NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017)) and six RAG pipelines. Table 1 shows the coverage when we prefetch $256$ clusters. From the table, we observe that IVF cluster overlap rates are consistently high across a range of datasets and pipelines. For instance, even the lowest reported values remain above 61.6% (for SubQ).
@@
Figure 6. (a) The overview of lookahead retrieval compared against CPU-offloaded retrieval, and (b) the system design of TeleRAG. After identifying clusters ($C_{\text{in}}$) for the initial query $q_{\text{in}}$, $C_{\text{in}}$ is transferred to the GPU while concurrently generating $q_{\text{out}}$. At retrieval time, the GPU searches prefetched clusters while the CPU handles remaining ones, before results are merged.
4. Design of TeleRAG
@@
Based on the high IVF cluster overlapping observation presented in §3.3, we propose TeleRAG, an efficient RAG inference system that utilizes lookahead retrieval to prefetch a minimal set of likely IVF clusters to GPU, and accelerate the retrieval process. In this section, we detail the design of lookahead retrieval and implementation of TeleRAG.
4.1. Overview
@@
Figure 6 presents the overview of TeleRAG with lookahead retrieval. In Figure 6, $q_{\text{in}}$ denotes the initial input to the pre-retrieval stage and $q_{\text{out}}$ denotes its output, i.e., the query given to the subsequent retrieval stage. We highlight the IVF clusters corresponding to $q_{\text{in}}$ and $q_{\text{out}}$ with a red background ($C_{\text{in}}$), and a blue background ($C_{\text{out}}$), respectively. Due to the semantic similarity of $q_{\text{in}}$ and $q_{\text{out}}$, there is significant overlap between $C_{\text{in}}$ and $C_{\text{out}}$, marked with a purple background ($C_{\text{overlap}}$).
Given this, the proposed lookahead retrieval technique in TeleRAG operates in three key steps:
  1. (1)
    Pre-retrieval & data transfer: During LLM generation, identify $C_{\text{in}}$ and transfer corresponding data to GPU memory, prioritizing clusters closest to $q_{\text{in}}$. This leverages GPU DMA for concurrent transfer of clusters data and compute of LLM generation.
  2. (2)
    GPU similarity search: After identifying $C_{\text{out}}$, the GPU efficiently performs a similarity search on the predicted clusters ($C_{\text{overlap}}$) already in its memory.
  3. (3)
    CPU similarity search: Concurrently with the GPU search, the CPU processes the similarity search for the missed clusters ($C_{\text{miss}}$) that were not prefetched to the GPU, reducing the CPU’s overall workload.
In summary, our lookahead retrieval speeds up retrieval by overlapping data prefetching to the GPU with LLM generation and performing concurrent similarity searches on both the GPU (for predicted clusters) and CPU (for the remainder), significantly reducing the CPU’s computational workload.
4.2. Finding the Optimal Prefetch Amount
@@
A key challenge in TeleRAG is to balance the benefit of reducing retrieval latency by prefetching data against the overhead of CPU-GPU transfers. Prefetching more clusters reduces the subsequent retrieval time but can also extend the transfer phase; if it extends beyond the LLM generation window, we lose the advantage of overlap and potentially introduce additional delay. Still, additional delays in data transfer may be worthwhile if they substantially reduce retrieval latency. To guide the choice of the optimal amount of data to prefetch, we develop a mathematical model based on profile information.
Mathematical model. Here, we denote $b_{p}$ as the number of bytes to prefetch and $B$ as the CPU-GPU bandwidth. The optimal amount of data to prefetch is denoted as $b_{p}^{*}$. To start, we let $t_{1}$ represent the combined time of prefetching and pre-retrieval LLM generation, and $t_{2}$ represents the retrieval time. We have: $t_{1}=\max(t_{\text{LLM}},t_{p})$ and $t_{2}=\max(t_{c},t_{g})$, where $t_{\text{LLM}}$ is the time for LLM generation, $t_{p}$ is prefetching time, $t_{c}$ is the CPU retrieval time, and $t_{g}$ is the GPU retrieval time. The objective is to minimize $t_{1}+t_{2}$.
@@
(1) $t_{1}=\begin{cases}t_{\text{LLM}},&\text{if }b_{p}\leq B\cdot t_{\text{LLM}},\\[6.0pt] \frac{b_{p}}{B},&\text{if }b_{p}>B\cdot t_{\text{LLM}},\end{cases}$
@@
As shown in §3.1, GPU retrieval ($t_{g}$) is generally much faster than CPU retrieval ($t_{c}$). Thus, we assume $t_{c}\gg t_{g}$. As CPU has a limited parallelism, $t_{c}$ usually grows proportionally with the number of clusters to be processed (Jiang et al., 2023a):
(2) $t_{2}=t_{c}=r_{\mathrm{miss}}\times n_{\mathrm{probe}}\times t_{cc},$
@@
    (4) $\frac{\mathrm{d}}{\mathrm{d}b_{p}}(t_{1}+t_{2})=0\implies\frac{1}{B}+\frac{\mathrm{d}r_{\mathrm{miss}}}{\mathrm{d}b_{p}}\cdot n_{\mathrm{probe}}\cdot t_{cc}=0.$
@@
    (5) $b_{p}^{*}=B\times n_{\mathrm{probe}}\times t_{cc}\times\Delta r_{\mathrm{miss}},$
@@
4.3. Design Details
Sorting Optimization. TeleRAG accelerates the final $k$-argmin sorting step of IVF search by leveraging GPU. Because each cluster may contain thousands of points, sorting is particularly well-suited to GPU parallelism (Arkhipov et al., 2017). To enable this, TeleRAG transfers the scalar distance values computed on the CPU for $C_{\text{miss}}$ to the GPU. Unlike full vector data, this transfer incurs negligible overhead. The GPU then performs a global sort over the combined distances from $C_{\text{overlap}}$ and $C_{\text{miss}}$, delivering end-to-end speedups than the CPU-only retrieval systems.
Prefetch Target. TeleRAG’s prefetching mechanism targets a specific number of bytes ($b_{p}$) rather than the number of clusters to ensure more predictable performance. Since IVF cluster sizes are often highly uneven, targeting a fixed cluster count can lead to unstable data loading times, whereas a byte limit provides a clearer upper bound on the transfer duration based on bandwidth.
When selecting clusters to meet the prefetch byte budget ($b_{p}$), the system adds whole clusters sequentially based on proximity to the query. If adding the next closest cluster would exceed the budget, it is omitted entirely rather than being partially loaded, ensuring a clean division where each cluster is processed exclusively by either GPU or CPU.
Multi-Round Support. For multi-round RAG involving the same input query, the system performs a full prefetch (up to the budget) only in the first round, leveraging high cluster similarity across rounds. In subsequent rounds, it incrementally fetches only the additional required clusters that were not loaded previously, optimizing data transfer while respecting the memory budget.
Batch Support. For batch inference, a challenge for lookahead retrieval is that each query requires a different set of IVF clusters. In our design, we fix the prefetching budget to the amount we got from §4.2 irrespective of batch size, and equally distribute the budget to each query in a batch. As a result, although the average cluster hit rate per request will decrease for larger batch sizes, we can ensure each query’s retrieval search time is equally accelerated, forming a balanced batched retrieval.
Multi-GPU support. For multi-GPU systems, we perform a greedy search based on embedding similarities of queries to group similar queries into mini-batches, and distribute the mini-batches among GPUs. The greedy search is efficient and incurs negligible overhead: In our profiling for up to the batch size of 256, the latency to perform greedy search is less than $0.1$ s, which is minimal compared to the end-to-end latency of each batch.
4.4. Implementation Details
TeleRAG is implemented in Python and also leverages PyTorch’s (Ansel et al., 2024) operator ecosystem for efficient computation. The datastore index is initially constructed using FAISS (Douze et al., 2024), and its data structures, such as IVF centroids and cluster data, are converted to PyTorch tensors.